OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@E.ali

OpenStudy (phi):

can you come up with another way to write x^2 ?

OpenStudy (anonymous):

I can't use x * x

OpenStudy (phi):

you could do a*b= x^2 x*x works, but if a= x^3 what is b ?

OpenStudy (phi):

you could also do a square root of x^4

OpenStudy (anonymous):

would b = x^-1?

OpenStudy (anonymous):

What is your mean ?

OpenStudy (phi):

yes, so x^-1 * x^3 = x^2 you could write x^-1 as 1/x so x^3/x = x^2

OpenStudy (phi):

you could do fractions. let a = x^(1/2) what is b so that a*b= x^2

OpenStudy (anonymous):

b = 3/2

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

OH! OK ! Did u get it ? Need my help ?

OpenStudy (phi):

because x^(1/2) is sqrt(x) you can write x^(1/2) * x^(3/2) as \[\sqrt{x} \cdot x^{\frac{3}{2}} = x^2 \]

OpenStudy (anonymous):

I think phi got most of it, thanks though ali

OpenStudy (anonymous):

You re welcome :

OpenStudy (phi):

which looks like a nice complicated way to write x squared

OpenStudy (anonymous):

haha, thanks

OpenStudy (anonymous):

Do u have any question ?

OpenStudy (anonymous):

No, I got them answered